Christin Demin invested part of her P240,000 in a firm that pays 12% per annum while the rest in a company that pays 7% per annum. If by the end of a year the two investments yielded a total interest of P21,600 how much was invested at each rate?
Answer by stanbon(75887)

Christin Demin invested part of her P240,000 in a firm that pays 12% per annum while the rest in a company that pays 7% per annum. If by the end of a year the two investments yielded a total interest of P21,600 how much was invested at each rate?
-------
Equations:
Quantity:: t+s = 240,000
Interest::12t+7s = 2160000
-----------------------------------
Modify for elimination::
7t + 7s = 7*240000
12t + 7s = 2160000
-----------------------
Subtract and solve for "t"::
5t = 480000
---
t = 96000 (amt. invested at 12%)
s = 240000-96000 = 144000 (amt invested at 7%)
-----
Cheers,
Stan H.
-----------